# How Internal Structure Shapes the Metallicity of Giant Exoplanets

### v1

@@ Abstract
Context. The composition and internal structure of gas giant exoplanets encode key information about their formation and subsequent evolution.
Aims. We investigate how different interior structure assumptions affect the inferred bulk metallicity and its correlation with planetary mass.
Methods. For a sample of 44 giant exoplanets (0.12–5.98 $M_{\rm J}$), we compute evolutionary models with CEPAM and retrieve their bulk metallicities under three structural hypotheses: Core+Envelope (CE), Dilute Core (DC), and Fully Mixed (FM).
Results. Across all structures, we recover a significant positive correlation between total heavy-element mass ($M_{Z}$) and planetary mass ($M$), and a negative correlation between metallicity ($Z$) and $M$ (also for $Z/Z_{\star}$ vs. $M$). DC structures yield metallicities comparable to CE models, regardless of the assumed gradient extent. Increasing atmospheric metallicity raises the inferred bulk metallicity, as enhanced opacities slow planetary cooling. Non-adiabatic DC models can further increase the retrieved metallicity by up to 35%. Sensitivity analyses show that the mass–metallicity anti-correlation is primarily driven by low-mass, metal-rich planets ($M<0.2\,M_{\rm J}$), while massive planets ($\gtrsim 1\,M_{\rm J}$) exhibit unexpectedly high metallicities ($Z\sim 0.1$$0.3$).
Conclusions. Improved constraints on convective mixing, combined with upcoming precise measurements of planetary masses, radii, and atmospheric compositions from missions such as PLATO and Ariel, will enable more robust inferences of interior structures and formation pathways for gas giant planets.
@@ 1 Introduction
Inferring a planet’s bulk composition and internal structure provides a powerful diagnostic of its formation pathway (Helled & Bodenheimer, 2011; Mordasini et al., 2016; Turrini et al., 2021). Gas giants represent an ideal case study in this context: as the largest planets, they must have formed early, when ample gas and solids were still available for accretion. Their size and rapid growth mean that they preserve valuable information about the composition of the protoplanetary disk (Pollack et al., 1996; Helled & Morbidelli, 2021; Mordasini et al., 2016). Hence, constraining their composition is key to probing the conditions under which planets form.
@@
The observational landscape now includes over 6,0006{,}000 confirmed exoplanets (NASA Exoplanet Archive, 2025), with a large, well‐characterised subset of giants having measured masses and radii from space and ground-based surveys (e.g. Kepler, TESS, WASP, HAT/HATSouth). The large dataset allows not only inferences of heavy‐element content in individual planets, but also statistical analyses across populations, enabling the identification of universal trends constrained by bulk properties.
@@
Interior models aim to reproduce the observed bulk properties by assuming specific distributions of heavy elements. These models often explore two simple end-member configurations: a Core+Envelope structure, where a distinct heavy-element core lies beneath a H/He envelope, and a Fully Mixed structure, where the heavy elements are homogeneously distributed throughout the planet (Thorngren et al., 2016; Howard et al., 2025b). However, gravity-field measurements and ring seismology for Jupiter and Saturn suggest more complex interiors. Interior models tailored to match these data suggest the interior of the two gas giants have compositional gradient of heavy elements or ’dilute cores’ (Wahl et al., 2017; Debras & Chabrier, 2019; Mankovich & Fuller, 2021; Miguel et al., 2022; Militzer et al., 2022; Howard et al., 2023; Ziv et al., 2024). Nevertheless, efforts in applying such a dilute core structure to exoplanetary studies is still embryonic. Recent retrieval studies have begun testing such structural prescriptions in individual exoplanets. For instance, Bloot et al. (2023) applied homogeneous and inhomogeneous models within a Bayesian framework to a limited sample, highlighting strong degeneracies that prevent quantitative discrimination between the two. Similarly, van Dijk & Miguel (2025) extended this approach by incorporating hot Jupiters, showing that current observations allow only qualitative assessments for a few well-characterized cases. Motivated by these findings, we extend interior modeling beyond the traditional end-member cases to include composition gradients and dilute-core structures that capture gradual heavy-element distributions.
@@
Figure 1: Three main internal structures tested in this study, with their respective heavy elements structure profiles: Core+Envelope (CE): all the heavy elements situated in a well defined core, Dilute Core, a gradient applied to the planet’s core, and Fully Mixed: an homogeneous envelope throughout the planet.
@@
In this work, we assess how the inferred bulk metallicity ($Z$) depends on the assumed interior structure for a sample of gas giant exoplanets. We further examine how different structural assumptions affect the resulting population-level trends. This study is important for the interpretation of data from JWST and upcoming missions such as PLATO and Ariel, which will provide precise measurements of planetary radii, ages, and atmospheric compositions, enabling tighter constraints on exoplanet interiors and formation histories. (Rauer et al., 2014; Tinetti et al., 2018).
@@ 2.1 Model parameters
We used CEPAM (Guillot & Morel, 1995; Howard et al., 2024) to compute planetary evolution. CEPAM numerically solves the 1D planetary structure and thermal-evolution equations (mass conservation, hydrostatic equilibrium, energy transport, and energy conservation) coupled to equations of state, that set density and entropy as functions of P, T, and composition. The sample size consists of 44 exoplanets taken from the Otegi et al. (2020); Parc et al. (2024) with masses between 0.12 and 5.98 $M_{\rm J}$. As in Howard et al. (2025b), we considered exoplanets with radius, mass, age, and $T_{eq}$ measurements and with relative measurement uncertainties smaller than 5 and 10 % for radius and mass, respectively. Exoplanets with equilibrium temperatures $T_{eq}$ ¿ 1000 K were excluded from this study, since additional energy sources can significantly inflate their radii (Guillot et al., 2006; Thorngren & Fortney, 2018). Such inflation mechanisms are not considered, and only planets for which our models reproduce the observed radius at the reported age were included.
The ice-to-rock ratio is kept constant for all the tested structures, and assumed to be a 1:1 ratio. For hydrogen and helium, the ratio between the two elements is assumed to be protosolar ($Y/(X+Y)$ = 0.27) (Asplund et al., 2021).
As shown by Howard et al. (2025b), the equation of states (EoS) used are important, and they affect the results inferred by the evolution models. The EoSs used in this study are the SESAME water and dry sand for ices and rocks (Lyon & Johnson, 1992), and the CMS19+HG23 EoS for H-He (Chabrier et al., 2019; Howard & Guillot, 2023).
The atmosphere in the model is treated as a non-gray, fully radiative atmosphere. Opacities are treated as two bands (visible and IR), and the mean opacities have been tabulated accordingly (Freedman et al., 2008, 2014; Parmentier et al., 2015). Opacity enhancement caused by heavy elements follows the methodology described in Valencia et al. (2013). For simplicity, we do not account for compositional mixing of heavy elements (the structure does not evolve over time). For all structures, the P-T profile is assumed to be adiabatic in the planet’s interior.
@@ 2.2 Tested interior structures
We considered different interior structures for the planetary dataset. The three main tested structures are Core+Envelope, Fully Mixed and Dilute Core. They can be described as follows:
  • Core+Envelope (CE): this structure assumes a well-defined core consisting of heavy elements, with the envelope being composed of H-He. As the metallicity in the core $Z_{\rm core}$ is equal to 1, we use the core mass $M_{\rm core}$ as a free parameter to infer bulk metallicities. For the basic Core+Envelope scenario, no heavy elements are inputted in the envelope and the atmosphere ($Z_{\rm env}$ = $Z_{\rm atm}=0$), where $Z_{\rm env}$ is the metallicity in the envelope and $Z_{\rm atm}$ in the atmosphere).
  • Core+Envelope with stellar atmosphere (CEA): To assess the contribution of metals mixed into the envelope, we allow heavy elements in the upper planetary layers $Z_{\rm env}=Z_{\rm atm}>0$.
  • Fully Mixed (FM): this structure assumes homogeneous composition throughout the planetary interior. No core is present. The bulk metallicity is equal to $Z_{\rm env}=Z_{\rm atm}$.
  • Dilute Core (DC): this structure represents a gradual distribution of heavy elements. Similarly to what done in Miguel (2022); Bloot et al. (2023); Howard et al. (2023), the structure is being described with the following function:
    $Z(m)=Z_{\text{env}}+\frac{Z_{\text{dilute}}-Z_{\text{env}}}{2}\left[1-\operatorname{erf}\!\left(\frac{m-m_{\text{dilute}}}{\delta m}\right)\right],$ (1)
    where $Z_{\rm dilute}$ is the maximum heavy–element fraction in the dilute core, the parameter $\delta m$ controls the steepness of the gradient, and $m_{\rm dilute}$ is the mass coordinate at which the gradient is steepest (see Fig. 1). $\delta m$ has been chosen to be fixed at 0.125 (as variations in this parameter are not believed to be substantially affecting planetary Z estimate (Bloot et al., 2023)). For the basic Dilute Core structure, $Z_{\rm env}=Z_{\rm atm}=0$.
  • Dilute Cores with Stellar and (Super)stellar atmospheres (DCA and DCA3): similar to DC, but now we allow atmospheric metallicity. The atmospheric abundances tested are matching the host star’s ($Z_{\star}$) metal abundance (so $Z_{\rm atm}=Z_{\star}$). Similar to Jupiter’s atmospheric enrichment (Helled & Stevenson, 2024), a more enriched case with $Z_{\rm atm}=3\,\times Z_{\star}$ has also been tested (named DCA3). To note that in these cases $m_{\rm dilute}$ will be fixed to a value of 0.5.
For each structure consistent with the observational constraints from each planet, we report $M_{Z}$ (heavy-element mass) and $Z$ (bulk metallicity), related by $Z=M_{Z}/M$, $M$ being the planetary mass.
@@ 2.3 Root-find $Z$ retrieval
In this section, we describe the statistical tool designed to retrieve the possible ranges of $Z$ that explain each planet’s radius for every structural assumption. The retrieval of planetary bulk metallicities $Z$ was designed as an inversion of the evolution model: instead of using $M$ and $Z$ to predict $(R,\text{Age})$, we seek the metallicity that reproduces the observed $(M,R,\text{Age})$ within their uncertainties. Ourapproach incorporates observational uncertainties by treating the mass and radius as normally distributed around their observed values ($M\sim\mathcal{N}(M_{\text{obs}},\sigma_{M})$, $R\sim\mathcal{N}(R_{\text{obs}},\sigma_{R})$) , while the age is assumed to follow a uniform distribution ($\text{Age}\sim\mathcal{U}(\text{Age}_{\min},\text{Age}_{\max})$). The method proceeds in three steps:
@@
  • Step 1 – Evolution runs: We first generated a set of cooling tracks with CEPAM across a broad grid of input parameters. Metallicities were sampled in the range $0\leq Z\leq 1$ with steps of 0.025, while masses were varied around the observed mean value, including the central value, $\pm 1\sigma$, $\pm 2\sigma$, and $\pm 3\sigma$. This ensemble of runs ensures that the parameter space captures both the observational uncertainties in mass and the possible degeneracies between metallicity and mass in shaping planetary radii. A dense coverage of $Z$ values is particularly important for detecting any degeneracies in the MMZZRR relation.
@@
  • Step 2 – Grid interpolation: The model outputs from CEPAM (describable as a Radius/Age relation for different structures and chemical abundances) were interpolated to build a continuous parameter grid. While linear interpolation is generally sufficient, we employed cubic spline interpolation (via scipy.interpolate.interp1d) since it better resolves curved behaviours and possible degeneracies in the cooling tracks. The interpolated grid was refined to a resolution of $\Delta Z=0.0025$, corresponding to 400 metallicity steps [^1] The choice of step sizes reflects a balance between accuracy and computational time. Tests on two representative planets with different MMZZ step combinations showed that denser sampling in metallicity is more critical than in mass. Such steps are assumed to be precise enough for the whole dataset. Future data with better observational constrains, or simply applying such a grid to Jupiter and Saturn, might be too coarse and unreliable, and it would require adjustments.. This step provides smooth predictions of planetary radii across the three-dimensional space $(M,Z,\text{Age})$, which are later required for the inversion procedure.
  • Step 3 – Root–finding and sampling:
@@
    Figure 2: Retrieved curves for CE, FM, and DC0.25-0.5-0.75 structures across three different planetary cases for different masses. CE curves are shown in purple, FM in yellow, and the DC curves are in black, and can be distinguished through the different line style.
@@
    To invert the grid and estimate $Z$, we repeatedly sampled observational values of mass, radius, and age. For each planet, we drew $n=5000$ tuples $(M,R,Age)$. For each tuple, Brent’s method was applied to solve
@@
    where $R_{\text{pred}}$ is the interpolated model prediction and $R_{\text{obs}}$ is the sampled observational value. This procedure yields the metallicity $Z$ that reproduces the observed radius. Repeating this process 5000 times provides a posterior distribution of $Z$, capturing both measurement uncertainties and intrinsic model degeneracies. A similar approach can be found in Müller & Helled (2023).
The root-finding retrieval provides the posterior distributions of metallicity $Z$ for each assumed interior structure and for every exoplanet analysed. These distributions were smoothed using a Kernel Density Estimate (KDE), and the corresponding mean and standard deviation of $Z$ were extracted and subsequently used to investigate planetary trends.
Figure 3: Retrieved curves for CE, FM, Core+Envelope + Stellar atmospheric composition (CEA), Dilute Core with 1 and 3 $Z_{\star}$ (DCA and DCA3), in vinaccia and orange colour respectively. Dashed lines (in wine colour and orange colour) show the atmospheric abundance assumed for DCA and DCA3 models respectively. Such lines will act as a lower threshold for these models as the assumed $R_{\text{atm}}$ will be kept fixed for each planet case.

### v2

@@ Abstract
Context. The composition and internal structure of gas giant exoplanets encode key information about their formation and evolution.
Aims. We investigate how different assumed interior structures affect the inferred bulk metallicity and its correlation with planetary mass.
Methods. For a sample of 44 giant exoplanets (0.12–5.98 $M_{J}$), we computed evolutionary models with CEPAM and retrieved their bulk metallicities under three structural hypotheses: core+envelope (CE), dilute core (DC), and fully mixed (FM).
Results. Across all structures, we recover a significant positive correlation between total heavy-element mass ($M_{Z}$) and planetary mass ($M$), and a negative correlation between bulk metallicity ($Z$) and $M$ (also for $Z/Z_{\star}$ vs $M$). Dilute core structures yield metallicities comparable to CE models, regardless of the assumed extent of the composition gradient. Increasing atmospheric metallicity augments the inferred bulk metallicity, as enhanced opacities slow planetary cooling. Non-adiabatic DC models can further increase the retrieved metallicity by up to 35%. We find that the mass–metallicity anti-correlation is primarily driven by low-mass, metal-rich planets ($M<0.2\,M_{J}$), and that massive planets ($\gtrsim 1\,M_{J}$) can exhibit unexpectedly high metallicities ($Z\sim 0.1$$0.3$).
Conclusions. Improved constraints on convective mixing, combined with upcoming accurate measurements of planetary masses, radii, and atmospheric compositions from missions such as PLATO and Ariel, will provide further constraints on interior structure and formation models of gas giant planets.
@@ 1 Introduction
Inferring a planet’s bulk composition and internal structure provides a powerful diagnostic of its formation pathway (Helled and Bodenheimer, 2011; Mordasini et al., 2016; Turrini et al., 2021). Gas giant planets represent an ideal case study in this context: as the largest planets, they must have formed early, when ample gas and solids were still available for accretion. Their high masses and rapid growth mean that they preserve valuable information about the composition of the protoplanetary disk (Pollack et al., 1996; Helled and Morbidelli, 2021; Mordasini et al., 2016). Hence, constraining their composition is key to probing the conditions under which planets form.
@@
The observational landscape now includes over 6,0006{,}000 confirmed exoplanets (NASA Exoplanet Archive, 2025), with a large, well‐characterised subset of giants having measured masses and radii from space and ground-based surveys (e.g. Kepler, TESS, WASP, and HAT/HATSouth). The large dataset allows inferences of heavy‐element content in individual planets and statistical analyses across populations, enabling the identification of more general trends constrained by bulk properties.
@@
Interior models are designed to reproduce the observed bulk properties by assuming specific distributions of heavy elements. These models often explore two simple end-member configurations: a core+envelope (CE) structure, where a distinct heavy-element core is surrounded by a H/He envelope, and a fully mixed (FM) structure, where the heavy elements are homogeneously distributed throughout the planet (Thorngren et al., 2016; Howard et al., 2025b). However, gravity-field measurements and ring seismology for Jupiter and Saturn suggest more complex interiors. Interior models tailored to match these data suggest that the interiors of the two gas giants consist of compositional gradients of heavy elements or ’dilute (fuzzy) cores’ (Wahl et al., 2017; Debras and Chabrier, 2019; Mankovich and Fuller, 2021; Miguel et al., 2022; Militzer et al., 2022; Howard et al., 2023; Ziv et al., 2024). Nevertheless, efforts in applying such a fuzzy or dilute core (DC) structure to exoplanetary studies are still embryonic. Recent retrieval studies have begun testing such structural prescriptions in individual exoplanets. For instance, Bloot et al. (2023) applied homogeneous and inhomogeneous models within a Bayesian framework to a limited sample, highlighting strong degeneracies that prevent quantitative discrimination between the two. Similarly, van Dijk and Miguel (2025) extended this approach by incorporating hot Jupiters, showing that current observations allow only qualitative assessments for a few well-characterised cases. In this study, we extend interior modelling beyond the traditional end-member cases to include composition gradients and DC structures that capture gradual heavy-element distributions.
@@
Figure 1: Three internal structures tested in this study, with their respective heavy element structure profiles: CE, all the heavy elements situated in a well-defined core; DC, a gradient applied to the planet’s core; and FM: a homogeneous envelope throughout the planet.
@@
In this work, we assess how the inferred bulk metallicity ($Z$) depends on the assumed interior structure for a sample of gas giant exoplanets. We further examine how different structural assumptions affect the resulting population-level trends. This study is important for the interpretation of JWST data (Carone et al., 2025; Deka et al., 2025) as well as future observations of giant planet atmospheres. Atmospheric measurements are currently being used to further constrain the planetary interiors (Thorngren and Fortney, 2019; Wilkinson et al., 2024; Knierim and Helled, 2025). Future missions such as PLATO and Ariel will provide accurate measurements of planetary radii, masses, ages, and atmospheric compositions, allowing tighter constraints on exoplanet interiors and formation histories. (Rauer et al., 2014; Tinetti et al., 2018).
@@ 2.1 Model parameters
We used the Code d’Evolution Planetaire Adaptif et Modulaire (CEPAM) (Guillot and Morel, 1995; Howard et al., 2024) to simulate the planetary evolution. CEPAM numerically solves the 1D planetary structure and thermal-evolution equations (mass conservation, hydrostatic equilibrium, energy transport, and energy conservation) coupled to EoSs, which set density and entropy as functions of $P$, $T$, and composition. The initial catalogue of giant planets from Otegi et al. (2020); Parc et al. (2024) was first restricted to those with all four required observational parameters (mass, radius, age, and $T_{\rm eq}$) available. This requirement alone substantially reduced the sample size. From this subset, we applied additional filtering, requiring relative uncertainties smaller than 5% in radius and 10% in mass, following Howard et al. (2025b), which yelded a final sample of 44 planets with masses between 0.12 and 5.98 $M_{\rm J}$. Exoplanets with equilibrium temperatures $T_{\rm eq}$ ¿ 1000 K were excluded from this study, since additional energy sources can significantly inflate their radii (Guillot et al., 2006; Thorngren and Fortney, 2018).
The ice-to-rock ratio was kept constant for all the tested structures and was assumed to be a 1:1 ratio. For hydrogen and helium, the ratio between the two elements was assumed to be protosolar ($Y/(X+Y)$ = 0.27) (Asplund et al., 2021).
The equations of state (EoSs) used are important and are known to affect the inferred composition (Baraffe et al., 2008; Miguel et al., 2016; Müller et al., 2020; Howard et al., 2025b). The EoSs used in this study are the SESAME water and dry sand for ices and rocks (Lyon and Johnson, 1992), and the CMS19+HG23 EoS for H-He (Chabrier et al., 2019; Howard and Guillot, 2023).
The atmosphere was treated as a non-grey, fully radiative atmosphere. Opacities were treated as two bands (visible and IR), and the mean opacities were tabulated accordingly (Freedman et al., 2008, 2014; Parmentier et al., 2015). Opacity enhancement caused by heavy elements follows the methodology described in Valencia et al. (2013).
For all internal structures, the $P$-$T$ profile was assumed to be adiabatic, and we did not consider compositional mixing of heavy elements. The initial entropy was not prescribed directly but instead was determined by the assumed initial planetary radius. The initial radii are typically in the range of 1–1.5 $R_{J}$, depending on the planetary mass. The choice of the initial radius is not expected to significantly influence the evolution when considering sufficiently old planets and adiabatic cooling (Marley et al., 2007; Müller and Helled, 2023). For completeness, we also considered a few cases where we assumed a non-adiabatic interior (see Section 4.1).
@@ 2.2 Tested interior structures
We considered different interior structures for the planetary dataset. The three main tested structures include the following:
  • Core+envelope: This structure assumes a well-defined core consisting of heavy elements, with the envelope being composed of H-He. As the metallicity in the core $Z_{\rm core}$ is unity, we used the core mass $M_{\rm core}$ as a free parameter to infer bulk metallicities. For the basic CE scenario, there are no heavy elements in the envelope: $Z_{\rm env}$ = $Z_{\rm atm}=0$), where $Z_{\rm env}$ is the metallicity in the envelope and $Z_{\rm atm}$ in the atmosphere.
  • Core+envelope with stellar-abundance atmosphere (CEA): To assess the contribution of metals mixed into the envelope, we allowed heavy elements in the upper planetary layers $Z_{\rm env}=Z_{\rm atm}>0$.
  • Fully mixed: This structure assumes a homogeneous composition throughout the planetary interior. No core is present. The bulk metallicity equals to $Z_{\rm env}=Z_{\rm atm}$.
  • Dilute core: This structure represents a gradual distribution of heavy elements. Following Miguel (2022); Bloot et al. (2023); Howard et al. (2023), we describe the structure using the following function:
    $Z(m)=Z_{\rm env}+\frac{Z_{\rm dilute}-Z_{\rm env}}{2}\left[1-\operatorname{erf}\!\left(\frac{m-m_{\rm dilute}}{\delta m}\right)\right],$ (1)
    where $Z_{\rm dilute}$ is the maximum heavy–element fraction in the DC, the parameter $\delta m$ controls the steepness of the gradient, and $m_{\rm dilute}$ is the mass coordinate at which the gradient is steepest (see Fig. 1). The value $\delta m$ was fixed at 0.125 (as variations in this parameter are not believed to substantially affect planetary Z estimates (Bloot et al., 2023)). In the basic DC structure, $Z_{\rm env}=Z_{\rm atm}=0$.
  • Dilute cores with stellar- and super-stellar abundance atmospheres (DCA and DCA3): This structure is the same as that for DC but allows for a non-zero atmospheric metallicity. The atmospheric abundances tested match the host star’s ($Z_{\star}$) metal abundance ($Z_{\rm atm}=Z_{\star}$). Similar to Jupiter’s atmospheric enrichment (Helled and Stevenson, 2024), a more enriched case with $Z_{\rm atm}=3\,\times Z_{\star}$ has also been tested (DCA3). Note that in these cases, $m_{\rm dilute}$ was fixed to a value of 0.5.
For each structure that is consistent with the observational constraints for a given planet, we report $M_{Z}$ (heavy-element mass) and $Z$ (bulk metallicity), related by $Z=M_{Z}/M$, where $M$ is the planetary mass.
@@ 2.3 Root-find $Z$ retrieval
In this section, we describe the statistical tool designed to retrieve the possible ranges of $Z$ that reproduce each planetary radius for every structural assumption. The retrieval of planetary bulk metallicities $Z$ was designed as an inversion of the evolution model: instead of using $M$ and $Z$ to predict $(R,\text{Age})$, we seek the metallicity that reproduces the observed $(M,R,\text{Age})$ within their uncertainties. Our approach incorporates observational uncertainties by treating the mass and radius as normally distributed around their observed values ($M\sim\mathcal{N}(M_{\text{obs}},\sigma_{M})$ and ($R\sim\mathcal{N}(R_{\text{obs}},\sigma_{R})$) , while the age is assumed to follow a uniform distribution ($\text{Age}\sim\mathcal{U}(\text{Age}_{\min},\text{Age}_{\max})$). The atmospheric metallicity was not used as an additional constraint, since this measurement is not available for the majority of the planets in our sample. The method proceeds in three steps:
@@
  • Step one – Evolution runs: We first generated a set of cooling tracks with CEPAM across a broad grid of input parameters. Metallicities were sampled in the range $0\leq Z\leq 1$ with steps of 0.025, while masses were varied around the observed mean value, and sampled at the central value, $\pm 1\sigma$, $\pm 2\sigma$, and $\pm 3\sigma$. This ensemble of runs ensures that the parameter space captures both the observational uncertainties in mass and the possible degeneracies between metallicity and mass in shaping planetary radii. A dense coverage of $Z$ values is particularly important for detecting degeneracies in the MMZZRR relation.
@@
  • Step two – Grid interpolation: The model outputs from CEPAM (describable as a radius/age relation for different structures and chemical abundances) were interpolated to build a continuous parameter grid. While linear interpolation is generally sufficient, we employed cubic spline interpolation (via scipy.interpolate.interp1d (Virtanen et al., 2020)) since it better resolves curved behaviours and possible degeneracies in the cooling tracks. The interpolated grid was refined to a resolution of $\Delta Z=0.0025$, corresponding to 400 metallicity steps [^1] The choice of step sizes reflects a balance between accuracy and computational time. Tests on two representative planets with different MMZZ step combinations showed that denser sampling in metallicity is more critical than in mass. These steps are sufficiently small for our dataset. Future data with better observational constraints, or simply applying such a grid to Jupiter and Saturn, might render this sampling too coarse and unreliable and would require adjustments.. This step provides smooth predictions of planetary radii across the 3D space $(M,Z,\text{Age})$, which are later required for the inversion procedure.
  • Step three – Root–finding and sampling:
@@
    Figure 2: Retrieved curves for CE, FM, and DC0.25-0.5-0.75 structures (varying $m_{\rm dilute}$ values) across three different planetary cases for different masses. The CE curves are shown in purple, FM in yellow, and DC in black, and can be distinguished through the different line style.
@@
    To invert the grid and estimate $Z$, we repeatedly sampled observational values of mass, radius, and age. For each planet, we drew $n=5000$ tuples $(M,R,\text{Age})$. For each tuple, Brent’s method was applied to solve:
@@
    where $R_{\text{pred}}$ is the interpolated model prediction and $R_{\text{obs}}$ is the sampled observational value. This procedure yields the metallicity $Z$ that reproduces the observed radius. Repeating this process $n$ times provides a posterior distribution for a prior of $Z\sim\mathcal{U}(0,1)$, capturing both measurement uncertainties and intrinsic model degeneracies. A similar approach was implemented in Müller and Helled (2023).
Root-finding retrieval provides the posterior distributions of metallicity $Z$ for each assumed interior structure and for every exoplanet analysed. These distributions were smoothed using a kernel density estimate (KDE), and the corresponding mean and standard deviation of $Z$ were extracted and used to investigate planetary trends.
Figure 3: Retrieved curves for CE, FM, CEA, DCA, and DCA3. The dashed lines (in burgundy and orange) show the atmospheric abundances assumed for DCA and DCA3 models, respectively. Such lines act as a lower threshold for these models as the assumed $Z_{\rm atm}$ was kept fixed for a given planet.